Calculate the net charge on a substance consisting
of a combination of 7.2 × 1013 protons
and 9.0×1013 electrons. The elemental charge
is 1.6 × 10−19 C.
Answer in units of C.
1
Expert's answer
2018-02-08T07:47:07-0500
The total charge
Q=N_1 e+N_2 (-e)=(N_1-N_2 )e
Q=(9.0-7.2)×〖10〗^13×1.6×〖10〗^(-19)=2.88×〖10〗^(-6) C=2.88 μC.